OpenStudy (anonymous):

The vertex of the parabola y=x^2-7x+9 is at

OpenStudy (anonymous):

what you want to do with the equation? find the x?

OpenStudy (anonymous):

this is bthe question

OpenStudy (anonymous):

the*

OpenStudy (anonymous):

oh okay. factorise it as like this y= a(x-b)^2 + c. you know how to do it right?

OpenStudy (anonymous):

no i dont i kinda of forgot cud u write it out i may know how to if i see it written out as a problem

OpenStudy (anonymous):

wait. let me work it out first

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

well this is what my teacher taught me y= x^2 - 7x +9 y= (x^2 - 7x + __ ) + 9 y= (x^2 -7x + 12.5) + 9 - 12.5 y= (x-3.5)^2 - 3.25 ---> the answer can you figure it out?

OpenStudy (anonymous):

is the 2nd step missing something

OpenStudy (anonymous):

yeah. because we dont know what's the thing, so we leave it blank. you know how to do it?